. Gears Inc. issued $10 million par value of bonds in 20X7 with an annual coupon rate of 8.0%. The bonds were issued at a price of $1,050 per $1000 par value, and they mature in ten years. The total interest cost of these bonds over their life is:
A. $8,500,000.
B. $8,400,000.
C. $8,000,000.
D. $7,500,000
Answer #1

Answer:

Correct answer is:

D. $7,500,000

Explanation:

Par value = $10,000,000

Annual coupon amount =10000000 * 8% = $800,000

Issue price = $1050

Value of issues = 10000000 * 1050/1000 = $10,500,000

Maturity = 10 years

To calculate yield%, we will use RATE function of excel:

= RATE (nper, pmt, pv, fv,type)

= RATE (10, 800000, -10500000, 10000000, 0)

= 7.27891%

Below is the bond amortization schedule:

As calculated above:

Total interest cost = $7,500,000

As such option D is correct and other options A, B and C are incorrect.